OpenStudy (anonymous):

There is a 30 percent chance of rain today and a 60 percent chance of rain tomorrow. If the two events are independent, what is the probability that it will rain today or tomorrow?

OpenStudy (anonymous):

How many possible outcomes do we have for these two events together?

OpenStudy (anonymous):

30/100 60/100

OpenStudy (anonymous):

Close, but I mean how many possible ways could it rain or not rain over the next two days?

OpenStudy (anonymous):

if the question was today and tomorrow the prob. would be .3 * .6 = .18 or 18 %

OpenStudy (anonymous):

The question's actually ambiguous - it could refer to having rain on either today or tomorrow (but not both), or at least one of the next two days; this changes the answer.

OpenStudy (anonymous):

P(A or B) = P(A) + P(B)

OpenStudy (anonymous):

an 'and' related to independent events are multiplied together. so today AND tomorrow = .3 * .6 but today OR tomorrow - are they added? - that would make 90% - seems a bit high!

OpenStudy (anonymous):

But we are looking for P( A & B) = P(A) * P(B)

OpenStudy (anonymous):

That's true, sorry these are unrelated.

OpenStudy (anonymous):

If we're looking for the exclusive or (one of the two days raining), then we could just subtract the probabilities of both days raining or not raining.

OpenStudy (anonymous):

Otherwise the probability will eventually exceed 100 ;)

OpenStudy (anonymous):

Correct you want to find 1- P(~A AND ~B)

OpenStudy (anonymous):

In that case, P(Day 1 Raining and Day 2 Raining) = .6 * .3 = .18

OpenStudy (anonymous):

P(Day 1 Not Raining and Day 2 Not Raining) = .4 * .7 = .28

OpenStudy (anonymous):

1 - P(Both Raining - P (Both Not Raining) = 1 - .18 - .28 = .54

OpenStudy (anonymous):

hmmm - my statistics knowledge is very rusty

OpenStudy (anonymous):

I don't think they care if it rains both days or just one of the two.

OpenStudy (anonymous):

In the other case, then P = 1 - P(Both Not Raining) = 1 - .28 = .72

OpenStudy (anonymous):

The probability that it will rain today or tomorrow is 1-.28 = .72

OpenStudy (anonymous):

yes - makes sense